# Identification and Optimal Nonlinear Control of Turbojet Engine Using Koopman Eigenfunction Model

### v2

@@ 6.4.2 Disturbance rejection
Figure 26: Disturbances of (a) altitude and (b) Mach number
Figure 27: Disturbance rejection results
7 Conclusions
To the author’s best knowledge, based on a rather extensive literature review, this is the first paper utilizing the KOT for the design of an optimal GTE controller. Furthermore, it is also first to combine the SINDy algorithm with the KOT for estimation of GTE governing equation with a subsequent Koopman canonical transform to the KEM. Another contribution is the constrained optimization of a low-order set of eigenvalues using metaheuristic global optimization tools, e.g. GA and PSO, and the projection of the autonomous trajectories in the time domain, providing an optimal set of eigenvalues for the eigenfunction construction from data. The objective function rejects the unstable, too oscillatory, and too lightly damped eigenvalues. Moreover, the algorithm can switch from distinct real and complex eigenvalues to repeated ones when the distance of any two eigenvalues is smaller then a specified margin. The discovered set of eigenfunctions and eigenvalues can also be extended by exploitation of the algebraic structure. However, by a product of two eigenfunctions we obtain a new eigenfunction that always decays faster than the original two functions and thus it less contributes to the dynamics.
The KEM with the Kalman observer was capable of an accurate state prediction even outside the region of validity of the LPV system. Furthermore, it allows for more aggressive integral control balanced by targeting of individual modes, where the modes with the lowest damping are the most important ones. This is captured by the reduced overshoots compared to PI and LPV-PI controllers. Furthermore, the LPV-PI and IMC controllers both struggle with the large deceleration, which points at too small PI gains in the low-RPM region for the deceleration. However, further increase of these gains would cause larger overshoots during small accelerations for the LPV-PI controller and increased sensitivity to noise for the IMC controller. Thus, the locally linearized controllers are overall effectively outperformed by the optimized KEM LQGI.
Both the identification method and the subsequently designed controller proved to be efficient for control of a single-spool turbojet GTE. For multi-spool turbojet and turbofan GTEs, the principle of temporal projection onto exponential trajectories is challenging and poses limitations. The problem is primarily caused by the ambiguity in coefficients K encoding both the unknown initial conditions of eigenfunctions and the mode amplitudes. These cannot be easily decoupled like in the case of 1st-order systems, where the variation of K for different trajectories is caused only by the changing initial conditions of $\Phi$. This will be a subject of future research. Fortunately, the optimization of eigenvalues is the only problem as the presented eigenfunction identification procedure works in higher dimensions as well.
In the future research, we will extend this framework to multi-rotor and turbofan engines.
@@ 7 Conclusions
Author contribution The author confirms sole responsibility for the following: study conception and design, simulation data collection, analysis and interpretation of results, and manuscript preparation.
@@
Funding The author declares that this research was supported by institutional funding provided by government of Czech Republic.
[…]
Data availability Data will be made available on request.
@@
[…]
Ethics approval and consent to participate Not applicable.
[…]
Consent for publication Not applicable.
[…]
Materials availability Not applicable.
@@ Appendix A Parameter-affine LPV approximation of the input dynamics
The input dynamics given by $\nabla\mathbf{\Phi}(N)g(N)$ are relatively complex as each of the entries is given by a product of two linear combinations of nonlinear basis functions which, in the case of $\nabla\mathbf{\Phi}$, are different for each entry. Thus, it would be suitable to approximate this complicated expression by an affine LPV representation and, possibly, even polytopic form.
@@
$\nabla\mathbf{\Phi}(N)g(N)=\sum_{i=1}^{m}\rho_{i}(N)\mathbf{\tilde{G}_{i}}=\mathbf{\tilde{G}}\boldsymbol{\rho},$ (99)
where $\mathbf{\tilde{G}}=\begin{bmatrix}\mathbf{\tilde{G}_{1}}&\mathbf{\tilde{G}_{2}}&\cdots&\mathbf{\tilde{G}_{m}}\end{bmatrix}$ are basis matrices of the decomposition and $\boldsymbol{\rho}=\begin{bmatrix}\rho_{1}&\rho_{2}&\cdots&\rho_{m}\end{bmatrix}^{\textbf{T}}$ are nonlinear scalar scheduling functions.
The primary goal is to optimize $\boldsymbol{\rho}$, parameterized by p, solving the following minimization problem:
p $=\arg\min_{\textbf{p}}\|\nabla\mathbf{\Phi}(N)g(N)-\mathbf{\tilde{G}}\boldsymbol{\rho}(N,\textbf{p})\|_{F}^{2},$
$\mathbf{\tilde{G}}$ $=\left(\boldsymbol{\rho\rho}^{\textbf{T}}-\alpha_{g}\textbf{I}\right)^{-1}\mathbf{\boldsymbol{\rho}G^{T}}\hskip 5.69054pt\bigg{|}_{\mathbf{p}=\text{const.}},$ (100)
where p is the vector of nonlinear parameters of the scheduling functions, $\alpha_{g}$ is the regularization parameter and $\textbf{G}=\nabla\mathbf{\Phi}(N)g(N)$.
@@
$\mathbf{L}=e^{\mathbf{\Lambda}\Delta t},\quad\mathbf{G_{d}=\Lambda^{-1}(L-I)\tilde{G}\boldsymbol{\rho}=\tilde{G}_{d}\boldsymbol{\rho}},$ (101)
@@
This problem can be handled with known optimization solvers, e.g. GD, GA or PSO. The latter two are capable of a quick search in the parametric space, whereas the GD can be utilized to analytically fine-tune the solution. However, the linear coefficients would need to be optimized as well leading to high-dimensional gradients and slower overall performance. Thus the meta-heuristic global optimization tools might be preferred.
The results of LPV decomposition with KEM prediction MAE, using different number of LFs, are depicted in Fig. 28. It can be concluded that the approximation is sufficient for $m=8$.
Figure 28: Comparison of affine LPV fitting for different number of LFs as basis functions
The computation of the input dynamics in each time step reduces only to an evaluation of 8 nonlinear functions and further matrix operations, which are rather fast. This is an improvement compared to the evaluation of $6\times 10$ functions for the gradient and 5 functions for $g(N)$. The model can be also converted to the polytopic form.

### v6

@@ Acknowledgements
Author contribution The author confirms sole responsibility for the following: study conception and design, simulation and data collection, analysis and interpretation of results, and manuscript preparation.
@@
Funding The author declares that this research was supported by funding provided by the government of the Czech Republic.
@@
Data availability Data will be made available on reasonable request.
@@ Appendix A Parameter-affine LPV approximation of the input dynamics
The input dynamics given by $\nabla\mathbf{\Phi}(N)g(N)$ are relatively complex as each of the entries is given by a product of two linear combinations of nonlinear basis functions, which, in the case of $\nabla\mathbf{\Phi}$, are even different for each entry. Thus, it would be suitable to approximate this complicated expression by an affine LPV representation, accelerating the computation of the Koopman observer equations for real-time applications.
@@
$\nabla\mathbf{\Phi}(N)g(N)=\sum_{k=1}^{m}\rho_{k}(N)\mathbf{\tilde{G}_{k}}=\mathbf{\tilde{G}}\boldsymbol{\rho},$ (103)
where $\mathbf{\tilde{G}}=\begin{bmatrix}\mathbf{\tilde{G}_{1}}&\mathbf{\tilde{G}_{2}}&\cdots&\mathbf{\tilde{G}_{m}}\end{bmatrix}$ are basis matrices of the decomposition and $\boldsymbol{\rho}=\begin{bmatrix}\rho_{1}&\rho_{2}&\cdots&\rho_{m}\end{bmatrix}^{\intercal}$ are nonlinear scalar scheduling functions.
Let us denote $\textbf{G}=\nabla\mathbf{\Phi}(N)g(N)$. The primary goal is to optimize $\boldsymbol{\rho}$, parameterized by p, solving the following minimization problem:
p $=\arg\min_{\textbf{p}}\|\textbf{G}-\mathbf{\tilde{G}}\boldsymbol{\rho}(N,\textbf{p})\|_{F}^{2}\ ,$
$\mathbf{\tilde{G}}$ $=\big(\boldsymbol{\rho}(N)\boldsymbol{\rho}(N)^{\intercal}-\alpha_{g}\textbf{I}\big)^{-1}\mathbf{\boldsymbol{\rho}}(N)\mathbf{G^{T}}\hskip 5.69054pt\bigg|_{\mathbf{p}=\text{const.}}\ ,$ (104)
where p is the vector of nonlinear parameters of the scheduling functions and $\alpha_{g}$ is the regularization parameter.
@@
$\mathbf{L}$ $=e^{\mathbf{\Lambda}\Delta t},$ (105)
$\mathbf{G_{d}}$ $=\mathbf{\Lambda^{-1}(L-I)\tilde{G}\boldsymbol{\rho}}$
$=\mathbf{\tilde{G}_{d}\boldsymbol{\rho}},$ (106)
@@
This problem can be handled with known optimization solvers, e.g., GD, GA, or PSO. The latter two are capable of a quick search in the parametric space, whereas the GD can be utilized to analytically fine-tune the solution. However, the linear coefficients would need to be optimized as well, leading to high-dimensional gradients and slower overall performance. Thus, the meta-heuristic global optimization tools might be preferred.
The results of LPV decomposition with KEM prediction MAE, using different numbers of LFs, are depicted in Fig. 29. It can be concluded that the approximation is sufficient for $m=8$.
Figure 29: Comparison of affine LPV fitting for different numbers of LFs as basis functions
The computation of the input dynamics in each time step reduces only to an evaluation of 8 nonlinear functions and further matrix operations, which are rather fast. This is an improvement compared to the evaluation of $6\times 10$ functions for the gradient and 5 functions for $g(N)$. The model can also be converted to the polytopic form.